OpenStudy (anonymous):

If Jose secures a $4000 loan that he will pay back in three years with a monthly interest rate of 1%, what is his monthly payment?

OpenStudy (anonymous):

what is the formula that i use?

jimthompson5910 (jim_thompson5910):

\[\Large M = \frac{L*i}{1-(1+i)^{-n}}\] M = monthly payment L = amount loaned out i = monthly interest rate n = number of months (convert years to months) So, M = unknown for now L = 4000 i = 0.01 (convert percent form to decimal form) n = 36 (36 months = 3 years)